OpenStudy (anonymous):

The total cost for 8 bracelets, including shipping was $54. The shipping charge was $6. Determine the cost for each bracelet. Show your work

OpenStudy (anonymous):

total cost =54 shipping charge=6 cost of bracelets=54-6=48 $48 for 8 bracelets => $48/8 for each bracelet =$6

OpenStudy (anonymous):

yes ...., for each bracelet is $6